A hollow glass tube A thin-walled hollow circular
Answer #1

surface charge density of glass tube is,

\sigma = \frac{Q}{2 \pi RL}

now electric field at point P is given by,

The component dEx of that electric field along the direction of the axis perpendicular to the plane of the ring is,

dEx = dE cos theta.jpg

dE_{x} = dE \left ( \frac{b}{r} \right)

dE_{x} = k\frac{Q}{r^2} \left ( \frac{b}{r} \right)

whrer r is ,

r = \left [ \left ( \frac{R}{2} \right )^2 + (\omega -x)^2 \right ]^{\frac{1}{2}}

and b= (\omega -x)

hence,

E_{x} = k \int_{0}^{L} \frac{ \omega -x}{[(\frac{R}{2})^2+ (\omega -x)^2]^{\frac{3}{2}}} \left ( \frac{Q}{2 \pi RL} \right ) 2 \pi R dx

E_{x} = \frac{kQ}{L} \int_{0}^{L} \frac{ \omega -x}{[(\frac{R}{2})^2+ (\omega -x)^2]^{\frac{3}{2}}} dx

by integrating,

E_{x} = \frac{kQ}{L} \left [ \frac{2}{ \sqrt{[ 4(x- \omega )^2 + R^2 ]}} \right ] _{0}^{L}

E_{x} = \frac{kQ}{L} \left [ \frac{2}{ \sqrt{[ 4(L- \omega )^2 + R^2 ]}} - \frac{2}{ \sqrt{[ 4(- \omega )^2 + R^2]}} \right ]

by pluging the given values, of L , R and \omega

E_{x} = \frac{(8.99 \times 10^{9} Nm^2/C^2)(9 \times 10^{-9}C)}{0.05m} \left [ \frac{2}{ \sqrt{[ 4(0.05m- 0.1m )^2 + (0.03m)^2 ]}} - \frac{2}{ \sqrt{[ 4(- 0.1m )^2 + (0.03m)^2]}} \right ]

E_{x} = 8990Nm/C \left [ \frac{2}{ \sqrt{ 0.0109}} - \frac{2}{ \sqrt{0.0409}} \right ]\frac{1}{m}

E_{x} = 8990 \left [ 19.15- 9.88 \right ] \frac{N}{C}

E_{x} = 83313.02 \, \frac{N}{C}